OpenStudy (anonymous):

Ned currently has an account balance of $3,634.51. He opened the account 13 years ago with a deposit of $2,564.65. If the interest compounds twice a year, what is the interest rate on the account?

OpenStudy (anonymous):

no idea to do this some pneone please help immediately

OpenStudy (anonymous):

this web page explains it nicely http://www.purplemath.com/modules/expofcns4.htm

OpenStudy (anonymous):

man no one cant't help me

OpenStudy (mathstudent55):

The formula for future value is: F = P(1 + r/n)^(nt) Where F = future value, P = present value (or principal), r = interest rate written as a decimal, n = number of times interest is compounded per year, t = time in years

OpenStudy (anonymous):

Thanks

OpenStudy (mathstudent55):

You know F = 3634.51, t = 13, n = 2, P = $2564.65. Substitute all those values in the formula, and solve for r.

OpenStudy (anonymous):

what' the rate

OpenStudy (mathstudent55):

Solve this equation for t: 3634.51 = 2564.65*(1 + t/2)^(2*13)

OpenStudy (anonymous):

i got 1.47

OpenStudy (anonymous):

is that right :)

OpenStudy (mathstudent55):

BTW, I meant to write r where t is in the equation with numbers. The equation is still correct, but where I wrote t, replace it with r. Solve this equation for r. 3634.51 = 2564.65*(1 + r/2)^(2*13) r can't be 1.47. Since r is the interest rate as a decimal that would mean the interest rate was 147%. That's not right.

OpenStudy (anonymous):

the answer is 1.4

OpenStudy (anonymous):

percent

OpenStudy (mathstudent55):

3634.51 = 2564.65*(1 + r/2)^(2*13) First, divide both sides by 2564.65 3634.51/2564.65 = (1 + r/2)^26 After doing the division on the left side, take the 26th root of both sides.

OpenStudy (mathstudent55):

I don't have a calculator on me, so I can't check on the answer.

OpenStudy (anonymous):

what did u get as yur final answer

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

I don't have a calculator on me, so I can't check on the answer.

OpenStudy (mathstudent55):

I can tell you what to do step by step, and you use your calculator to get the answer.

OpenStudy (mathstudent55):

3634.51/2564.65 = (1 + r/2)^26 First, divide 3634.51 by 2564.65

OpenStudy (mathstudent55):

Ok, I found the Google calculator. I'm doing it now.

OpenStudy (mathstudent55):

I get r = 0.027

OpenStudy (anonymous):

ok i plugged it in and got 1.4%

OpenStudy (mathstudent55):

I got r = 0.027, so the interest rate is 2.7%

OpenStudy (anonymous):

ohh ok isee how u did it :0